OpenStudy (anonymous):

In need of help again. stuck on this problem and I've already had 4 people try to help me. 11√49a^5/4a^3

OpenStudy (anonymous):

put your problem in at http://www.wolframalpha.com/?fp=right

OpenStudy (anonymous):

it didnt work

OpenStudy (campbell_st):

quick question is it 1. \[\sqrt{\frac{49a^5}{4a^3}}\] or 2. \[\frac{\sqrt{49a^5}}{4a^3}\]

OpenStudy (anonymous):

(77/4) a^2 ? hard to know where the square roots stop

OpenStudy (anonymous):

\[11\sqrt{\frac{ 49a^5 }{ 4a^3 }}\]

OpenStudy (anonymous):

the answers gonna come out to be \[\frac{ 77a }{ 2}\] I just don't know how haha

OpenStudy (campbell_st):

ok... so look at it this way... cancel the a's first so its \[\sqrt{\frac{49a^2 \times a^3}{4 \times a^3}}\] what cancels...from the numerator and denominator..?

OpenStudy (campbell_st):

oops forgot the 11 outside.. pick that up in a min

OpenStudy (anonymous):

uh a^3?

OpenStudy (campbell_st):

ok... thats great so you have and applying a fule about radicals it can be split to the square root of the factors.. \[11 \sqrt{\frac{49a^2}{4}} = \frac{11 \sqrt{49a^2}}{4}\] to take the square root of the numerator and denominator which gives \[\frac{11 \times 7a}{2} \] and hence your answer

OpenStudy (anonymous):

Thank you so much!!